A manufacturer claims that 10% of women using the "pill" suffer from side effects. The Federal Drug Administration (FDA) believes that the manufacturer's claim is too low and decides to test the manufacturer's claim at α = 5 %. A random sample of 900 women who use the manufacturer's pill shows that 103 suffer side effects.

The standard error for this test is ___
Answer #1

SE=\sqrt{\frac{p(1-p)}{n}}=\sqrt{\frac{0.10*0.90}{900}}=0.01